OpenStudy (anonymous):

find the value of b satisfying the description the line through the points (1,5,) and (-3,b) with slope -2/7 the value of b is?

OpenStudy (chaise):

You have points: (x,y) (1,5) (-3,b) Your slope is m = -2/7 The formula for slope is: m= (y-y)/(x-x) Substitute your values in. Solve for b -2/7 = b-5/-3-1 -2/7=b-5/-3-1 -2/7=b-5/-4 8/7=b-5 b=43/7 ~ 6.1